An electron trapped in a one-dimensional box of length L = 1 ร… [a typical atomic diameter] isย 

described by the normalized wave function

๐›น = โˆš2

๐ฟ

sin

๐‘›๐œ‹

๐‘ฅ๐ฟ

ย In the ground state, what is the probability that the electron lies between (a) 0.09ย 

ร… and 0.11 ร… (b) 0.0 ร…ย 

ย and 0.25 ร…?


1
Expert's answer
2021-05-07T06:41:28-0400

a)

"P=\int_{0.09}^{0.11}\psi ^2dx=\int_{0.09}^{0.11}\frac 2L\sin^2\frac {n \pi x}{L} dx=0.00383,"

b)

"P=\int_{0.09}^{0.11}\psi ^2dx=\int_{0}^{0.25}\frac 2L\sin^2\frac {n \pi x}{L} dx=0.09085."